# FEDONet : Fourier-Embedded DeepONet for Spectrally Accurate Operator Learning

### v2

@@ 3.7 Kuramoto–Sivashinsky Equation
We evaluate the performance of FEDONet on the one-dimensional Kuramoto–Sivashinsky (KS) equation, a canonical model that exhibits spatiotemporal chaos due to the combined effects of nonlinear advection, destabilizing diffusion, and stabilizing fourth-order hyperviscosity. The governing equation is given by:
@@
Figure 17: Median-performing KS trajectory. Relative $L^{2}$ error: FEDONet = 16.36%, DeepONet = 89.63%.
@@
Figure 18: Best-performing KS test sample. FEDONet retains the amplitude and frequency content of the chaotic trajectory, yielding a low relative $L^{2}$ error of $10.77\%$.
Figure 19: Worst-performing KS test sample. Even in failure cases, FEDONet remains qualitatively reasonable. Relative $L^{2}$ error: FEDONet = $58.02\%$, versus near-complete divergence for DeepONet.
@@
Figure 20: KS Training convergence comparison. FEDONet achieves faster and more stable optimization due to implicit spectral regularization.
@@ 4 Summary
In this work, we propose the Fourier-Embedded DeepONet (FEDONet) as a principled extension to the classical DeepONet architecture for learning nonlinear operators governed by partial differential equations (PDEs). By injecting randomized Fourier features into the trunk input, FEDONet expands the hypothesis space of DeepONets, offering enhanced spectral representation, improved numerical conditioning, and faster convergence during training all without increasing runtime complexity. FEDONet consistently outperforms the vanilla DeepONet. Notably, it delivers dramatic reductions in relative $\ell_{2}$ error, most prominently in multiscale and high-frequency scenarios such as the Kuramoto-Sivashinsky equation (76% $\rightarrow$ 21%).
Table 1: Relative $\ell_{2}$ error (%) for each benchmark.
@@
2-D Poisson (elliptic) 5.80 2.32
1-D Burgers (conv./diff.) 4.01 2.63
Lorenz–63 (chaotic ODE) 2.03 0.46
Lorenz–96 (chaotic ODE) 23.40 22.90
Allen–Cahn (phase–field) 10.61 5.50
Kuramoto–Sivashinsky (spatio-chaos) 75.11 20.62
@@ 5 Future Work
The Fourier-Embedded DeepONet (FEDONet) architecture presents a compelling foundation for enhancing operator learning, yet several promising extensions remain unexplored. A natural next step is to generalize beyond sinusoidal embeddings by investigating alternative basis functions in the trunk network. Wavelets, Chebyshev polynomials, or orthogonal polynomials tailored to specific PDE families could offer improved localization or boundary adaptivity, especially for systems exhibiting sharp interfaces or multi-scale dynamics. Additionally, hybridizing spectral embeddings with structured architectures such as graph-based trunks or encoder–decoder networks may further boost performance on irregular domains or hierarchical solution spaces.
@@ Appendix A Whitening Effect of Fourier Feature Embeddings
Let $\phi(\zeta)=\sqrt{2}\,[\sin(2\pi B\zeta),\cos(2\pi B\zeta)]\in\mathbb{R}^{2M}$ be the Fourier feature embedding of an input $\zeta\in\mathbb{R}^{d}$, where each row of the matrix $B\in\mathbb{R}^{M\times d}$ is sampled i.i.d. from a Gaussian distribution:
$B_{ij}\sim\mathcal{N}(0,\sigma^{2}).$ (32)
@@
To see this, consider a single frequency row vector $B_{i}\in\mathbb{R}^{1\times d}$ drawn from $B$. The corresponding $2$-dimensional block of the embedding is:
$\phi_{i}(\zeta)=\sqrt{2}\begin{bmatrix}\sin(2\pi B_{i}\cdot\zeta)\\ \cos(2\pi B_{i}\cdot\zeta)\end{bmatrix}$. (34)
We now compute the second-order statistics of this embedding block under the assumption that $B_{i}\cdot\zeta$ is uniformly distributed over $[0,1]$ (a reasonable approximation due to the randomness in $B_{i}$ and the uniformity of $\zeta$). Using standard trigonometric integrals, we obtain:
$\mathbb{E}_{\zeta}[\sin^{2}(2\pi B_{i}\cdot\zeta)]$ $\approx\frac{1}{2},$ (35)
$\mathbb{E}_{\zeta}[\cos^{2}(2\pi B_{i}\cdot\zeta)]$ $\approx\frac{1}{2},$ (36)
$\mathbb{E}_{\zeta}[\sin(2\pi B_{i}\cdot\zeta)\cos(2\pi B_{i}\cdot\zeta)]$ $\approx 0.$ (37)
@@
$\mathbb{E}_{\zeta}\left[\begin{bmatrix}\sin(2\pi B_{i}\cdot\zeta)\\ \cos(2\pi B_{i}\cdot\zeta)\end{bmatrix}\begin{bmatrix}\sin(2\pi B_{i}\cdot\zeta)&\cos(2\pi B_{i}\cdot\zeta)\end{bmatrix}\right]\approx\frac{1}{2}I_{2}.$ (38)
@@
Since the rows of $B$ are sampled independently and the embeddings for each $B_{i}$ are uncorrelated, the full embedding vector $\phi(\zeta)\in\mathbb{R}^{2M}$ satisfies:
@@ Appendix B Theoretical Justification for Fourier Embedding Superset Property
$\phi(\zeta)=\left[\sin(2\pi B\zeta),\cos(2\pi B\zeta)\right],\quad B\in\mathbb{R}^{M\times d},\;B_{ij}\sim\mathcal{N}(0,\sigma^{2}).$ (41)
@@
Part II: Strict Containment. Consider the function $f(\zeta)=\sin(100\pi\zeta)$ over $\zeta\in[0,1]$. Due to the spectral bias of neural networks [49], this high-frequency function is difficult to approximate using shallow networks on raw input. However, if the frequency matrix $B$ used in the embedding $\phi(\zeta)$ includes rows close to $100$, then:
$\sin(2\pi B_{j}\zeta)\approx\sin(100\pi\zeta),$
@@ Appendix C Spectral Interpretation of Fourier-Embedded DeepONets
    The operator learning framework parallels well-established numerical discretization techniques, such as Petrov–Galerkin and spectral methods, but operates in a fully data-driven regime.

### v3

@@ 3.6 Kuramoto–Sivashinsky Equation
We evaluate the performance of FEDONet on the one-dimensional Kuramoto-Sivashinsky (KS) equation, a canonical model that exhibits spatiotemporal chaos due to the combined effects of nonlinear advection, destabilizing diffusion, and stabilizing fourth-order hyperviscosity. The governing equation is given by:
@@
Figure 15: Solving a parametric KS equation: Median performing test-sample. Relative $L^{2}$ error: FEDONet = 16.36%, DeepONet = 89.63%.
@@
Figure 16: Solving a parametric KS equation: Best performing test sample. FEDONet retains the amplitude and frequency content of the chaotic trajectory, yielding a low relative $L^{2}$ error of $10.77\%$.
Figure 17: Solving a parametric KS equation: Worst performing test sample. Even in failure cases, FEDONet remains qualitatively reasonable. Relative $L^{2}$ error: FEDONet = $58.02\%$, versus near-complete divergence for DeepONet.
@@
Figure 18: KS Training convergence comparison. FEDONet achieves faster and more stable optimization.
@@ 4 Summary
In this work, we propose the Fourier-Embedded DeepONet (FEDONet) as a principled extension to the classical DeepONet architecture for learning nonlinear operators governed by partial differential equations. By injecting randomized Fourier features into the trunk input, FEDONet expands the hypothesis space of DeepONets, offering enhanced spectral representation, improved numerical conditioning, and faster convergence during training all without increasing runtime complexity. FEDONet consistently outperforms the vanilla DeepONet. Notably, it delivers dramatic reductions in relative $L^{2}$ error, most prominently in multiscale and high-frequency scenarios such as the Kuramoto-Sivashinsky equation (76% $\rightarrow$ 21%).
Table 1: Relative $L^{2}$ error (%) for each benchmark.
@@
2-D Poisson (elliptic) 1.41 1.09
1-D Burgers (conv./diff.) 6.43 4.86
Lorenz-63 (chaotic ODE) 2.03 0.46
Allen-Cahn (phase–field) 10.61 5.50
Kuramoto-Sivashinsky (spatio-chaos) 75.11 20.62
@@ 5 Future Work
The Fourier-Embedded DeepONet (FEDONet) architecture presents a compelling foundation for enhancing operator learning, yet several promising extensions remain unexplored. A natural next step is to generalize beyond sinusoidal embeddings by investigating alternative basis functions in the trunk network. Wavelets, Chebyshev polynomials, or orthogonal polynomials tailored to specific PDE families could offer improved localization or boundary adaptivity, especially for systems exhibiting sharp interfaces or multi-scale dynamics. Additionally, hybridizing spectral embeddings with structured architectures such as graph-based trunks or encoder-decoder networks may further boost performance on irregular domains or hierarchical solution spaces.
@@ Acknowledgements
Declaration of Generative AI use
During the preparation of this work, the authors used ChatGPT and Grammarly to improve readability and language. After using these tools, the authors reviewed and edited the content as needed and take full responsibility for the content of the publication.
@@ Appendix A Whitening Effect of Fourier Feature Embeddings
Let $\phi(\zeta)=\sqrt{2}\,[\sin(2\pi Z\zeta),\cos(2\pi Z\zeta)]\in\mathbb{R}^{2M}$ be the Fourier feature embedding of an input $\zeta\in\mathbb{R}^{d}$, where each row of the matrix $Z\in\mathbb{R}^{M\times d}$ is sampled i.i.d. from a Gaussian distribution:
$Z_{ij}\sim\mathcal{N}(0,\sigma^{2}).$ (30)
@@
To see this, consider a single frequency row vector $Z_{i}\in\mathbb{R}^{1\times d}$ drawn from $Z$. The corresponding $2$-dimensional block of the embedding is:
$\phi_{i}(\zeta)=\sqrt{2}\begin{bmatrix}\sin(2\pi Z_{i}\cdot\zeta)\\ \cos(2\pi Z_{i}\cdot\zeta)\end{bmatrix}$. (32)
We now compute the second-order statistics of this embedding block under the assumption that $Z_{i}\cdot\zeta$ is uniformly distributed over $[0,1]$ (a reasonable approximation due to the randomness in $Z_{i}$ and the uniformity of $\zeta$). Using standard trigonometric integrals, we obtain:
$\mathbb{E}_{\zeta}[\sin^{2}(2\pi Z_{i}\cdot\zeta)]$ $\approx\frac{1}{2},$ (33)
$\mathbb{E}_{\zeta}[\cos^{2}(2\pi Z_{i}\cdot\zeta)]$ $\approx\frac{1}{2},$ (34)
$\mathbb{E}_{\zeta}[\sin(2\pi Z_{i}\cdot\zeta)\cos(2\pi Z_{i}\cdot\zeta)]$ $\approx 0.$ (35)
@@
$\mathbb{E}_{\zeta}\left[\begin{bmatrix}\sin(2\pi Z_{i}\cdot\zeta)\\ \cos(2\pi Z_{i}\cdot\zeta)\end{bmatrix}\begin{bmatrix}\sin(2\pi Z_{i}\cdot\zeta)&\cos(2\pi Z_{i}\cdot\zeta)\end{bmatrix}\right]\approx\frac{1}{2}I_{2}.$ (36)
@@
Since the rows of $Z$ are sampled independently and the embeddings for each $Z_{i}$ are uncorrelated, the full embedding vector $\phi(\zeta)\in\mathbb{R}^{2M}$ satisfies:
@@ Appendix B Theoretical Justification for Fourier Embedding Superset Property
$\phi(\zeta)=\left[\sin(2\pi Z\zeta),\cos(2\pi Z\zeta)\right],\quad Z\in\mathbb{R}^{M\times d},\;Z_{ij}\sim\mathcal{N}(0,\sigma^{2}).$ (39)
@@
Part II: Strict Containment. Consider the function $f(\zeta)=\sin(100\pi\zeta)$ over $\zeta\in[0,1]$. Due to the spectral bias of neural networks [51], this high-frequency function is difficult to approximate using shallow networks on raw input. However, if the frequency matrix $Z$ used in the embedding $\phi(\zeta)$ includes rows close to $100$, then:
$\sin(2\pi Z_{j}\zeta)\approx\sin(100\pi\zeta),$
@@ Appendix C Spectral Interpretation of Fourier-Embedded DeepONets
    The operator learning framework parallels well-established numerical discretization techniques, such as Petrov-Galerkin and spectral methods, but operates in a fully data-driven regime.